This is a trig question about degrees and radians that I don't understand.
A centrifuge is a machine that is used to seperate fluids. In a high school lab, the centrifuge has an angular speed of 400pi radians per second.
A. Find the rotation in revolutions per minute
B. What is the linear speed of a substance that is 8 centimeters away from the center?

Answer by lwsshak3(11628)

A centrifuge is a machine that is used to seperate fluids. In a high school lab, the centrifuge has an angular speed of 400pi radians per second.
***
A. Find the rotation in revolutions per minute.
400 π radians/sec*revolution/2π radians*60 sec/min
π, sec, radians cancel out, leaving revolution/min
400 π radians/sec*revolution/2π radians*60 sec/min=400*60/2=12000 rev/min
..
B. What is the linear speed of a substance that is 8 centimeters away from the center?
circumference=2πr=2π*8=16π cm
for each revolution the subtance travels 16π cm
12000 rev/min*16π/rev=192000π cm/min